### v2

@@ 5.1 Local linear convergence
$=-\sqrt{\bar{H}}P_{k}^{\top}(P_{k}H_{k}P_{k}^{\top}+\eta_{k}I_{s})^{-1}P_{k}H_{k}(x_{k}-\bar{x})+\sqrt{\bar{H}}P_{k}^{\top}(P_{k}H_{k}P_{k}^{\top}+\eta_{k}I_{s})^{-1}P_{k}(g_{k}-H_{k}(x_{k}-\bar{x}))+\sqrt{\bar{H}}(x_{k}-\bar{x})$
$=-A+B+\sqrt{\bar{H}}(x_{k}-\bar{x}),$ (53)
where $A:=\sqrt{\bar{H}}P_{k}^{\top}(P_{k}H_{k}P_{k}^{\top}+\eta_{k}I_{s})^{-1}P_{k}H_{k}(x_{k}-\bar{x})$ and $B:=\sqrt{\bar{H}}P_{k}^{\top}(P_{k}H_{k}P_{k}^{\top}+\eta_{k}I_{s})^{-1}P_{k}(g_{k}-H_{k}(x_{k}-\bar{x}))$. First let us bound $B$. In order to do so, we bound $\|P_{k}^{\top}(P_{k}H_{k}P_{k}^{\top}+\eta_{k}I_{s})^{-1}P_{k}\|$. Notice that by Assumption 5.1$(i)$ and Proposition 5.1, $P_{k}H_{k}P_{k}^{\top}\succ 0$ holds for $n\geq n_{0}$. Hence, since $\eta_{k}\geq 0$,
$\|P_{k}^{\top}(P_{k}H_{k}P_{k}^{\top}+\eta_{k}I_{s})^{-1}P_{k}\|\leq\|P_{k}^{\top}(P_{k}H_{k}P_{k}^{\top})^{-1}P_{k}\|.$
Hence we deduce that with probability at least $1-6\exp(-s)$,
$\|P_{k}^{\top}(P_{k}H_{k}P_{k}^{\top}+\eta_{k}I_{s})^{-1}P_{k}\|\leq\frac{\|P_{k}^{\top}\|^{2}}{\frac{(1-\varepsilon_{0})^{2}n}{2s}\sigma\bar{\lambda}}$ (54)
holds. Therefore, by Lemma 2.2, we have
$\|P_{k}^{\top}(P_{k}H_{k}P_{k}^{\top}+\eta_{k}I_{s})^{-1}P_{k}\|\leq\frac{2\mathcal{\bar{C}}}{{(1-\varepsilon_{0})^{2}}\sigma\bar{\lambda}}.$ (55)
By Taylor expansion at $\bar{x}$ of $\nabla f$, as in (50), and by subtracting $H_{k}(x_{k}-\bar{x})$ to both side, we obtain by Assumption 5 that
[…]
Hence, by (55) and (56), there exists a constant $\beta_{1}>0$ such that
$B\leq\|\sqrt{\bar{H}}\|\|P_{k}^{\top}(P_{k}H_{k}P_{k}^{\top}+\eta_{k}I_{s})^{-1}P_{k}\|\|(g_{k}-H_{k}(x_{k}-\bar{x}))\|\leq\beta_{1}\|x_{k}-\bar{x}\|^{2}.$ (57)
@@
$=\sqrt{\bar{H}}P_{k}^{\top}(P_{k}\bar{H}P_{k}^{\top}+\eta_{k}I_{s})^{-1}P_{k}\bar{H}(x_{k}-\bar{x})+\sqrt{\bar{H}}P_{k}^{\top}((P_{k}H_{k}P_{k}^{\top}+\eta_{k}I_{s})^{-1}P_{k}H_{k}$
$~{}~{}~{}~{}~{}~{}-(P_{k}\bar{H}P_{k}^{\top}+\eta_{k}I_{s})^{-1}P_{k}\bar{H})(x_{k}-\bar{x}).$ (58)
Notice that by Assumption 5, we have that $\|({\bar{H}}-{{H_{k}}})\|$ tends to 00. Therefore, we deduce from (55) and (57) that
@@
Therefore by (5.17), (57) and (5.17), we deduce that
@@
$\|\sqrt{\bar{H}}(x_{k+1}-\bar{x})\|\leq\left\|\left(I_{n}-\sqrt{\bar{H}}P_{k}^{\top}(P_{k}\bar{H}P_{k}^{\top})^{-1}P_{k}\sqrt{\bar{H}}\right)\sqrt{\bar{H}}(x_{k}-\bar{x})\right\|+o(\|x_{k}-\bar{x}\|).$ (59)
@@
Then similarly to the proof of Proposition 5.5 and similarly to (5.6), we have that with probability at least $1-6(\exp(-s)+\exp(-\frac{\mathcal{C}_{0}}{4}s))$,
@@
Therefore, by (59), we have that
@@
By Lemma 5.14, we have $o(\|x_{k}-\bar{x}\|)=o(\|\sqrt{\bar{H}}(x_{k+1}-\bar{x})\|)$, hence we deduce that when $k$ is large enough,
@@
We complete the proof using (40).
@@ 5.2 Impossibility of local super-linear convergence in general
holds with high probability when $k$ is large enough. This will prove that the results obtained in the previous section are optimal when $f$ is locally strongly-convex. Indeed, by local strong-convexity of $f$ and Hessian Lipschitz continuity (i.e. Assumption 5), there exists $l_{2}\geq l_{1}>0$ such that for $k$ large enough,
@@
where $\mathcal{C}$ is the constant that appears in (26).
@@
where $t_{k}$ is the step-size and $\eta_{k}>0$ is a parameter that tends to 00 when $k$ tends to infinity.
Let us fix $k$. Using a Taylor expansion of $t\mapsto\nabla f(\bar{x}+t(x_{k+1}-\bar{x}))$ around 00, as in (50), we have that
[…]
where the second inequality holds for $k$ large enough under Assumption 5. Hence, for $k$ large enough and under Assumption 5:
$\|x_{k+1}-\bar{x}\|\geq\frac{1}{2\lambda_{\max}(\nabla^{2}f(\bar{x}))}\|\nabla f(x_{k+1})\|,$ (61)
where $\lambda_{\max}(\cdot)$ denotes the largest eigenvalue. Using a Taylor expansion of $\nabla f$ around $x_{k}$, we have that
[…]
@@
By Assumption 5, the Hessian is $L_{H}$-Lipschitz in $B_{H}$. Since $x_{k}$ and ${x}_{k}+t(x_{k+1}-x_{k})\in B_{H}$ for $k$ large enough, we have that
@@
Hence (61) leads to
$\|x_{k+1}-\bar{x}\|\geq\frac{1}{2\lambda_{\max}(\nabla^{2}f(\bar{x}))}\left(\|g_{k}+H_{k}(x_{k+1}-x_{k})\|-L_{H}\|x_{k+1}-x_{k}\|^{2}\right).$ (62)
@@
Assume that Assumptions 5 and 5.2 hold and that $f$ is strongly convex locally around $\bar{x}$. There exists a constant $\beta>0$ such that if $k$ is large enough, then with probability at least $1-2\exp(-\frac{\mathcal{C}_{0}}{4}s)-2\exp(-s)$, we have
@@
Proof. We have
@@
where $M_{k}$ is defined in (6). If $k$ is large enough, $H_{k}$ is invertible by strong convexity of $f$. Notice that $\|(I_{n}-t_{k}H_{k}P_{k}^{\top}M_{k}^{-1}P_{k})g_{k}\|=\|H_{k}(H_{k}^{-1}-t_{k}P_{k}^{\top}M_{k}^{-1}P_{k})g_{k}\|$. Hence since for any invertible matrix $A$ we have $\|Ax\|\geq\frac{\|x\|}{\|A^{-1}\|}$, we deduce that
@@
Hence, by (63), we deduce that
$\|(H_{k}^{-1}-t_{k}P_{k}^{\top}M_{k}^{-1}P_{k})g_{k}\|^{2}$ (64)
@@
Since $H^{-1}_{k}g_{k}=P_{k}^{\top}z_{1}+z_{2}$ with $P_{k}z_{2}=0$, we have that $P_{k}H^{-1}_{k}g_{k}=P_{k}P_{k}^{\top}z_{1}$. Hence
@@
By Lemma 2.1, we have that with probability at least $1-2\exp(-\frac{\mathcal{C}_{0}}{4}s)$ that $\|P_{k}H^{-1}_{k}g_{k}\|\leq 2\|H^{-1}_{k}g_{k}\|$. Furthermore, by writing the singular value decomposition, $U\Sigma V^{\top}$, of $P_{k}^{\top}$, we have that  $\|P_{k}^{\top}(P_{k}P_{k}^{\top})^{-1}\|=\|U\Sigma^{-1}V^{\top}\|=\frac{1}{\sigma_{\min}(P_{k}^{\top})}$. Since $\sigma_{\min}(P_{k}^{\top})\geq\sqrt{\frac{n}{s}}-\mathcal{C}$ holds with probability at least $1-2e^{-s}$ (we only consider the first equation of (26)), we deduce that
@@
Hence, from (64) we have
$\|(H_{k}^{-1}-t_{k}P_{k}^{\top}M_{k}^{-1}P_{k})g_{k}\|^{2}$ (65)
@@
where we used that $2ab\leq a^{2}+b^{2}$ in the last inequality, and that $\left(1-\frac{2}{\sqrt{\frac{n}{s}}-\mathcal{C}}\right)>0$ holds by Assumption 5.2. Hence, from (65) we proved that
@@
$\|g_{k}+H_{k}(x_{k+1}-x_{k})\|\geq\frac{\sqrt{1-\frac{2}{\sqrt{\frac{n}{s}}-\mathcal{C}}}}{\|H_{k}^{-1}\|}\|x_{k+1}-x_{k}\|.$ (66)
Considering $k$ large enough, as $x_{k}$ tends to $\bar{x}$, we can bound, using Assumption 5, $\frac{1}{\|H_{k}^{-1}\|}\geq\frac{1}{2\|{\bar{H}}^{-1}\|}$, where we recall that $\bar{H}=\nabla^{2}f(\bar{x})$, which ends the proof.
@@
Assume that Assumptions 5 and 5.2 hold and that $f$ is locally strongly convex around $\bar{x}$. There exists a constant $c>0$ such that for $k$ large enough,
@@
Proof. From (62) and Proposition 5.18 we deduce that with probability at least $1-2\exp(-\frac{\mathcal{C}_{0}}{4}s)-2\exp(-s)$, when $k$ is large enough
@@
$\|x_{k+1}-\bar{x}\|\geq\frac{\beta}{4\lambda_{\max}(\bar{H})}\|x_{k+1}-x_{k}\|.$
Now we notice that
$\|x_{k+1}-x_{k}\|=\|P_{k}^{\top}M_{k}^{-1}P_{k}g_{k}\|\geq\sigma_{\min}(P_{k}^{\top})\|M_{k}^{-1}\|\|P_{k}g_{k}\|.$ (67)
Using Lemma 2.1 and the bound (26) on $\sigma_{\min}(P_{k}^{\top})$, we have that
$\sigma_{\min}(P_{k}^{\top})\|M_{k}^{-1}\|\|P_{k}g_{k}\|\geq\left(\sqrt{\frac{n}{s}}-\mathcal{C}\right)\|M_{k}^{-1}\|\frac{1}{2}\|g_{k}\|.$ (68)
@@
$\|M_{k}^{-1}\|\geq\frac{1}{2\mathcal{C}\lambda_{\max}(\bar{H})\frac{n}{s}}.$ (69)
Hence, by (67),(68) and (69) we have that there exists a constant $\kappa_{2}>0$ such that
@@
By (50) we have that
[…]
Hence, since $f$ is assumed to be locally strongly convex, by Assumptions 5 we have that for $k$ large enough:
$\|g_{k}\|\geq\frac{\lambda_{\min}(\bar{H})}{2}\|\bar{x}-x_{k}\|.$
Hence we have for $k$ large enough that
@@
Assume that Assumptions 5 and 5.2 hold and that $f$ is locally strongly convex around $\bar{x}$. There exists a constant $c>0$ such that for $k$ large enough,
@@
where $\bar{c}=(1-2\exp(-\frac{\mathcal{C}_{0}}{4}s)-2\exp(-s))c$, and where the expectation is taken with respect to the random variables $P_{1},\cdots,P_{k}$.
@@
Proof. The proof is very similar to the proof of Corollary 5.12. Let us consider the random variable $\mathbb{E}[\|x_{k+1}-\bar{x}\|\ |\ P_{1},\cdots,P_{k-1}]$. Let =$\mathcal{E}=\{\|x_{k+1}-\bar{x}\|\geq\bar{c}\|x_{k}-\bar{x}\|\ |\ x_{k}\}$ be an event with respect to the random variable $P_{k}$. Using the fact that $\|x_{k+1}-\bar{x}\|\geq 0,$ we obtain that
$\mathbb{E}\left[\|x_{k+1}-\bar{x}\|\ |\ P_{1},\cdots,P_{k-1}\right]$ $=\mathbb{E}\left[\|x_{k+1}-\bar{x}\|\ |\ P_{1},\cdots,P_{k-1},{\mathcal{E}}\right]P(\mathcal{E})+\mathbb{E}\left[\|x_{k+1}-\bar{x}\|\ |\ P_{1},\cdots,P_{k-1},{\bar{\mathcal{E}}}\right](1-P(\mathcal{E}))$
@@
Taking the expectation with respect to $P_{1},\cdots,P_{k-1}$ leads to the result.

### v3

@@ 5.1 Local linear convergence
(58) $=-\sqrt{\bar{H}}P_{k}^{\top}(P_{k}H_{k}P_{k}^{\top}+\eta_{k}I_{s})^{-1}P_{k}H_{k}(x_{k}-\bar{x})+\sqrt{\bar{H}}P_{k}^{\top}(P_{k}H_{k}P_{k}^{\top}+\eta_{k}I_{s})^{-1}P_{k}(g_{k}-H_{k}(x_{k}-\bar{x}))$
$\ \ \ +\sqrt{\bar{H}}(x_{k}-\bar{x})$
(59) $=-A+B+\sqrt{\bar{H}}(x_{k}-\bar{x}),$
where $A:=\sqrt{\bar{H}}P_{k}^{\top}(P_{k}H_{k}P_{k}^{\top}+\eta_{k}I_{s})^{-1}P_{k}H_{k}(x_{k}-\bar{x})$ and $B:=\sqrt{\bar{H}}P_{k}^{\top}(P_{k}H_{k}P_{k}^{\top}+\eta_{k}I_{s})^{-1}P_{k}(g_{k}-H_{k}(x_{k}-\bar{x}))$. First let us bound $B$. In order to do so, we bound $\|P_{k}^{\top}(P_{k}H_{k}P_{k}^{\top}+\eta_{k}I_{s})^{-1}P_{k}\|$. Notice that from $P_{k}H_{k}P_{k}^{\top}\succ 0$, $\eta_{k}\geq 0$ and Proposition 5.1, we have
(60) $\|P_{k}^{\top}(P_{k}H_{k}P_{k}^{\top}+\eta_{k}I_{s})^{-1}P_{k}\|\leq\|P_{k}^{\top}(P_{k}H_{k}P_{k}^{\top})^{-1}P_{k}\|\leq{\frac{\|P_{k}^{\top}\|^{2}}{\frac{n}{8s}\sigma\bar{\lambda}}}$
with probability at least $1-6\exp(-s)$. Therefore, by Lemma 2.2, we have
(61) $\|P_{k}^{\top}(P_{k}H_{k}P_{k}^{\top}+\eta_{k}I_{s})^{-1}P_{k}\|\leq\frac{8\mathcal{\bar{C}}}{\sigma\bar{\lambda}}.$
By Taylor expansion at $\bar{x}$ of $\nabla f$, as in Eq. 54, and by subtracting $H_{k}(x_{k}-\bar{x})$ to both sides, we obtain by Section 5 that
(62) $\|g_{k}-H_{k}(x_{k}-\bar{x})\|\leq\int_{0}^{1}\|\nabla^{2}f(\bar{x}+t(x_{k}-\bar{x}))-\nabla^{2}f(\bar{x})\|\|x_{k}-\bar{x}\|dt=O(\|x_{k}-\bar{x}\|^{2}).$
Hence, by Eq. 60 and Eq. 62, there exists a constant $\beta_{1}>0$ such that
(63) $B\leq\|\sqrt{\bar{H}}\|\|P_{k}^{\top}(P_{k}H_{k}P_{k}^{\top}+\eta_{k}I_{s})^{-1}P_{k}\|\|(g_{k}-H_{k}(x_{k}-\bar{x}))\|\leq\beta_{1}\|x_{k}-\bar{x}\|^{2}.$
@@
(64) $=\sqrt{\bar{H}}P_{k}^{\top}(P_{k}\bar{H}P_{k}^{\top}+\eta_{k}I_{s})^{-1}P_{k}\bar{H}(x_{k}-\bar{x})+\sqrt{\bar{H}}P_{k}^{\top}((P_{k}H_{k}P_{k}^{\top}+\eta_{k}I_{s})^{-1}P_{k}(H_{k}-\bar{H})(x_{k}-\bar{x}).$
Notice that by Section 5, we have that $\|({\bar{H}}-{{H_{k}}})\|$ tends to 0. Therefore, we deduce from Eq. 60 and Eq. 63 that
@@
Therefore by 5.17, Eq. 63 and 5.17, we deduce that
@@
(65) $\|\sqrt{\bar{H}}(x_{k+1}-\bar{x})\|\leq\left\|\left(I_{n}-\sqrt{\bar{H}}P_{k}^{\top}(P_{k}\bar{H}P_{k}^{\top})^{-1}P_{k}\sqrt{\bar{H}}\right)\sqrt{\bar{H}}(x_{k}-\bar{x})\right\|+o(\|x_{k}-\bar{x}\|).$
@@
Then similarly to the proof of Proposition 5.5 and similarly to 5.6, we have that with probability at least $1-6(\exp(-s)+\exp(-\frac{\mathcal{C}_{0}}{4}s))$,
@@
Therefore, by Eq. 65, we have that
@@
By Lemma 5.14, we have $o(\|x_{k}-\bar{x}\|)=o(\|\sqrt{\bar{H}}(x_{k}-\bar{x})\|)$, hence we deduce that when $k$ is large enough,
@@
We complete the proof using Eq. 44.
@@ 5.2 Impossibility of local super-linear convergence in general
holds with high probability when $k$ is large enough. This will prove that the results obtained in the previous section are optimal when $f$ is locally strongly-convex. Indeed, by local strong-convexity of $f$ and Hessian Lipschitz continuity (i.e. Section 5), there exists $l_{2}\geq l_{1}>0$ such that for $k$ large enough,
@@
where $\mathcal{C}$ is the constant that appears in Eq. 29.
@@
where $t_{k}$ is the step-size and $\eta_{k}>0$ is a parameter that tends to 0 when $k$ tends to infinity.
Let us fix $k$. Using a Taylor expansion of $t\mapsto\nabla f(\bar{x}+t(x_{k+1}-\bar{x}))$ around 0, as in Eq. 54, we have that
(66) $\|\nabla f(x_{k+1})\|\leq\int_{0}^{1}\|\nabla^{2}f(\bar{x}+t(x_{k+1}-\bar{x}))\|\|x_{k+1}-\bar{x}\|dt\leq\int_{0}^{1}2\lambda_{\max}(\nabla^{2}f(\bar{x}))\|x_{k+1}-\bar{x}\|dt,$
where $\lambda_{\max}(\cdot)$ denotes the largest eigenvalue, and the second inequality holds for $k$ large enough under Section 5. Hence, for $k$ large enough and under Section 5,
(67) $\|x_{k+1}-\bar{x}\|\geq\frac{1}{2\lambda_{\max}(\nabla^{2}f(\bar{x}))}\|\nabla f(x_{k+1})\|$
holds. Using a Taylor expansion of $\nabla f$ around $x_{k}$, we have that
$\nabla f(x_{k+1})=\nabla f({x_{k}})+\int_{0}^{1}\nabla^{2}f({x_{k}}+t(x_{k+1}-{x_{k}}))(x_{k+1}-{x_{k}})dt.$
@@
By Section 5, the Hessian is $L_{H}$-Lipschitz in $B_{H}$. Since $x_{k}$ and ${x}_{k}+t(x_{k+1}-x_{k})\in B_{H}$ for $k$ large enough, we have that for $t\leq 1$,
@@
Hence Eq. 67 leads to
(68) $\|x_{k+1}-\bar{x}\|\geq\frac{1}{2\lambda_{\max}(\nabla^{2}f(\bar{x}))}\left(\|g_{k}+H_{k}(x_{k+1}-x_{k})\|-L_{H}\|x_{k+1}-x_{k}\|^{2}\right).$
@@
Assume that Section 5 and Section 5.2 hold and that $f$ is strongly convex locally around $\bar{x}$. There exists a constant $\beta>0$ such that if $k$ is large enough, then with probability at least $1-2\exp(-\frac{\mathcal{C}_{0}}{4}s)-2\exp(-s)$, we have
@@
Proof. Recalling the updated rule $x_{k+1}=x_{k}-t_{k}P_{k}^{\mathsf{T}}M_{k}^{-1}P_{k}g_{k}$ in Algorithm 1, we have
@@
where $M_{k}$ is defined in Eq. 6. If $k$ is large enough, $H_{k}$ is invertible by strong convexity of $f$. Notice that $\|(I_{n}-t_{k}H_{k}P_{k}^{\top}M_{k}^{-1}P_{k})g_{k}\|=\|H_{k}(H_{k}^{-1}-t_{k}P_{k}^{\top}M_{k}^{-1}P_{k})g_{k}\|$. Hence since for any invertible matrix $A$ we have $\|Ax\|\geq\frac{\|x\|}{\|A^{-1}\|}$, we deduce that
@@
Hence, by Eq. 69, we deduce that
(70) $\|(H_{k}^{-1}-t_{k}P_{k}^{\top}M_{k}^{-1}P_{k})g_{k}\|^{2}$
@@
Since $H^{-1}_{k}g_{k}=P_{k}^{\top}z_{1}+z_{2}$ with $P_{k}z_{2}=0$, we have that $P_{k}H^{-1}_{k}g_{k}=P_{k}P_{k}^{\top}z_{1}$. Which implies (since $P_{k}P_{k}^{\top}$ is invertible with probability $1$) that $z_{1}=(P_{k}P_{k}^{\top})^{-1}P_{k}H^{-1}_{k}g_{k}$. Hence
@@
By Lemma 2.1, we have that with probability at least $1-2\exp(-\frac{\mathcal{C}_{0}}{4}s)$ that $\|P_{k}H^{-1}_{k}g_{k}\|\leq 2\|H^{-1}_{k}g_{k}\|$. Furthermore, by writing the singular value decomposition, $U\Sigma V^{\top}$, of $P_{k}^{\top}$, we have that  $\|P_{k}^{\top}(P_{k}P_{k}^{\top})^{-1}\|=\|U\Sigma^{-1}V^{\top}\|=\frac{1}{\sigma_{\min}(P_{k}^{\top})}$. Since $\sigma_{\min}(P_{k}^{\top})\geq\sqrt{\frac{n}{s}}-\mathcal{C}$ holds with probability at least $1-2e^{-s}$ (we only consider the first equation of Eq. 29), we deduce that
@@
Hence, from Eq. 70 we have
(71) $\|(H_{k}^{-1}-t_{k}P_{k}^{\top}M_{k}^{-1}P_{k})g_{k}\|^{2}$
@@
where we used that $2ab\leq a^{2}+b^{2}$ in the last inequality, and that $\left(1-\frac{2}{\sqrt{\frac{n}{s}}-\mathcal{C}}\right)>0$ holds by Section 5.2. Hence, from (71) we proved that
@@
(72) $\|g_{k}+H_{k}(x_{k+1}-x_{k})\|\geq\frac{\sqrt{1-\frac{2}{\sqrt{\frac{n}{s}}-\mathcal{C}}}}{\|H_{k}^{-1}\|}\|x_{k+1}-x_{k}\|.$
Considering $k$ large enough, as $x_{k}$ tends to $\bar{x}$, we can bound, using Section 5, $\frac{1}{\|H_{k}^{-1}\|}\geq\frac{1}{2\|{\bar{H}}^{-1}\|}$, where we recall that $\bar{H}=\nabla^{2}f(\bar{x})$, which ends the proof.
@@
Assume that Section 5 and Section 5.2 hold and that $f$ is locally strongly convex around $\bar{x}$. There exists a constant $c>0$ such that for $k$ large enough,
@@
Proof. From Eq. 68 and Proposition 5.18 we deduce that with probability at least $1-2\exp(-\frac{\mathcal{C}_{0}}{4}s)-2\exp(-s)$, when $k$ is large enough
@@
(73) $\|x_{k+1}-\bar{x}\|\geq\frac{\beta}{4\lambda_{\max}(\bar{H})}\|x_{k+1}-x_{k}\|.$
Since $f$ is assumed to be strongly convex, for all $\alpha\in(0,1)$, as $g_{k}^{\mathsf{T}}d_{k}\leq 0$. Hence we have that $t_{k}=1$Now we notice that
(74) $\|x_{k+1}-x_{k}\|=t_{k}\|P_{k}^{\top}M_{k}^{-1}P_{k}g_{k}\|\geq t_{k}\sigma_{\min}(P_{k}^{\top})\|M_{k}^{-1}\|\|P_{k}g_{k}\|.$
Using Lemma 2.1 (with $\varepsilon=1/2$) and the bound Eq. 29 on $\sigma_{\min}(P_{k}^{\top})$, we have that
(75) $t_{k}\sigma_{\min}(P_{k}^{\top})\|M_{k}^{-1}\|\|P_{k}g_{k}\|\geq t_{k}\left(\sqrt{\frac{n}{s}}-\mathcal{C}\right)\|M_{k}^{-1}\|\frac{1}{2}\|g_{k}\|.$
@@
(76) $\|M_{k}^{-1}\|\geq\frac{1}{2\mathcal{C}\lambda_{\max}(\bar{H})\frac{n}{s}}.$
Hence, by Eqs. 73, 74, 75, and 76 we have that there exists a constant $\kappa_{2}>0$ such that
@@
By Eq. 54 we have that
$g_{k}=\bar{H}(x_{k}-\bar{x})+\int_{0}^{1}(\nabla^{2}f(\bar{x}+t(x_{k}-\bar{x}))-\bar{H})(x_{k}-\bar{x})dt.$
Hence, since $f$ is assumed to be locally strongly convex, by Section 5 we have that for $k$ large enough:
$\|g_{k}\|\geq\frac{\lambda_{\min}(\bar{H})}{2}\|x_{k}-\bar{x}\|.$
Using (4.12), we have
$f(x_{k})-f(x_{k}+t_{k}^{\prime}d_{k})+\alpha t_{k}^{\prime}g_{k}^{\mathsf{T}}d_{k}\geq\frac{\bar{\mathcal{C}}n}{2s}L_{H}{t_{k}^{\prime}}^{2}\left\|d_{k}\right\|\left(\frac{c_{2}s\left\|g_{k}\right\|^{\gamma}}{\bar{\mathcal{C}}L_{H}n\left\|d_{k}\right\|}-t_{k}^{\prime}\right)\left\|M_{k}^{-1}P_{k}g_{k}\right\|^{2},$
and since $f$ is assume to be strongly convex, $\frac{\|g_{k}\|^{\gamma}}{\|d_{k}\|}$ is in the order of $\mathcal{O}(\frac{1}{\|g_{k}\|^{1-\gamma}})$, hence $t_{k}$ is bounded below by some constant for $k$ large enough. Hence we have for $k$ large enough that
@@
Assume that Section 5 and Section 5.2 hold and that $f$ is locally strongly convex around $\bar{x}$. Then for $k$ large enough,
@@
where $\bar{c}=(1-2\exp(-\frac{\mathcal{C}_{0}}{4}s)-2\exp(-s))c$ ($c$ is the same constant as in Theorem 5.20), and where the expectation is taken with respect to the random variables $P_{0},\cdots,P_{k}$.
@@
Proof. The proof is very similar to the proof of Corollary 5.12. Let us consider the random variable $\mathbb{E}[\|x_{k+1}-\bar{x}\|\ |\ P_{0},\cdots,P_{k-1}]$. Let $\mathcal{E}=\{\|x_{k+1}-\bar{x}\|\geq\bar{c}\|x_{k}-\bar{x}\|\ |\ x_{k}\}$ be an event with respect to the random variable $P_{k}$. Using the fact that $\|x_{k+1}-\bar{x}\|\geq 0,$ we obtain that
$\mathbb{E}\left[\|x_{k+1}-\bar{x}\|\ |\ P_{0},\cdots,P_{k-1}\right]$ $=\mathbb{E}\left[\|x_{k+1}-\bar{x}\|\ |\ P_{0},\cdots,P_{k-1},{\mathcal{E}}\right]P(\mathcal{E})$
$\ \ \ +\mathbb{E}\left[\|x_{k+1}-\bar{x}\|\ |\ P_{0},\cdots,P_{k-1},{\bar{\mathcal{E}}}\right](1-P(\mathcal{E}))$
@@
Taking the expectation with respect to $P_{0},\cdots,P_{k-1}$ leads to the result.